Price of x is 12 and price of y is 8. income is $600 U(x, y)=x^0.4 y^0.6

set up lagurangian, write down first order conditions, solve the system of equation in first order condition to find the optimal x and y and \lambda. explain how you interpret the value of  \lambda. and then, find the marginal utility of good x when the consumer chooses the optimal bundle.

please solve this step by step.
Answer #1

= x 4y66 Px = 12 Py = 8 I 100 (Price of (Pose of y). (Income Budget Constroint: x. Px + 4. Py = I > 12x +84 600 Set up lagrinEguate eguaten 4 and 6. > on 64900 (39) 04 > 6) - 2025 -> X = 2:254 - 12 put (6 in (3). 12x + 8y - 600 12. (2.254) + 8y = 6010. • уеб de Mur du lux Mux = 04 X-Ow6yo: Nuo 106 * aut y = 17:14 and x = 38.56. MUX = 0.4 / 17:14 10.6 Mux .0.2459 38.56 Mar